# Benchmarking Vector, Graph and Hybrid Retrieval Augmented Generation (RAG) Pipelines for Open Radio Access Networks (ORAN)

### v1

@@ Abstract
Generative AI (GenAI) is expected to play a pivotal role in enabling autonomous optimization in future wireless networks. Within the ORAN architecture, Large Language Models (LLMs) can be specialized to generate xApps and rApps by leveraging specifications and API definitions from the RAN Intelligent Controller (RIC) platform. However, fine-tuning base LLMs for telecom-specific tasks remains expensive and resource-intensive. Retrieval-Augmented Generation (RAG) offers a practical alternative through in-context learning, enabling domain adaptation without full retraining. While traditional RAG systems rely on vector-based retrieval, emerging variants such as GraphRAG and Hybrid GraphRAG incorporate knowledge graphs or dual retrieval strategies to support multi-hop reasoning and improve factual grounding. Despite their promise, these methods lack systematic, metric-driven evaluations, particularly in high-stakes domains such as ORAN. In this study, we conduct a comparative evaluation of Vector RAG, GraphRAG, and Hybrid GraphRAG using ORAN specifications. We assess performance across varying question complexities using established generation metrics: faithfulness, answer relevance, context relevance, and factual correctness. Results show that both GraphRAG and Hybrid GraphRAG outperform traditional RAG. Hybrid GraphRAG improves factual correctness by 8%, while GraphRAG improves context relevance by 7%.
@@ II Related Work
Figure 2: Experimental pipeline, illustrating the key components and data flow across the system.
@@ III-D Dataset
For this study, we utilized a corpus of 74 documents from the ORAN Alliance Specifications [28]. The evaluation was conducted using the ORAN-Bench-13K dataset [14], a benchmark specifically designed to assess the performance of LLMs within the ORAN context. The dataset categorizes questions into three levels of complexity: Easy, Intermediate, and Hard. A total of 600 questions were selected for evaluation, ensuring coverage across a range of topics including network architecture and components, analytics and monitoring, anomaly detection, and protocol interpretation. Each question includes four answer options along with a ground truth label. The categorization reflects increasing levels of difficulty: (i) Easy: Questions targeting foundational concepts or factual knowledge (Simple QA) (ii) Intermediate: Questions requiring moderate reasoning, comprehension, or application of concepts (Complex Reasoning QA) (iii) Hard: Questions demanding deep understanding of ORAN standards and the ability to synthesize information across multiple documents (Multi-hop Reasoning)  [6].
@@ IV Evaluation
Figure 4: Responses generated by Vector RAG, GraphRAG, and Hybrid GraphRAG for a benchmark question from ORAN‑Bench‑13K
@@ IV-B Results
Figure 3 visualizes the performance of each model across four core evaluation metrics stratified by question difficulty. To provide a broader perspective, Table II summarizes the average metric scores for each model, aggregated across all difficulty levels. The results are discussed below, organized by evaluation metric.
In terms of faithfulness, all three models perform similarly on easy questions, each achieving an average score of approximately 0.77. However, on medium and hard questions, graph-based pipelines outperform vector-based RAG. Overall, Hybrid GraphRAG and GraphRAG (both scoring 0.59) outperform RAG (0.55), suggesting that their generated responses are more consistently grounded in the retrieved context and are less prone to hallucination.
TABLE II: Overall Average Performance Across All Question Types
@@
Faithfulness 0.59 0.59 0.55
Factual Correctness 0.50 0.58 0.48
Context Relevance 0.11 0.04 0.10
Answer Relevance 0.74 0.72 0.73
In terms of factual correctness, Hybrid GraphRAG achieves the highest scores across all difficulty levels, with an overall accuracy of 0.58. This is attributed to its ability to compensate for the limitations of individual retrieval methods, leveraging vector-based retrieval when graph-derived context is insufficient, and vice versa. GraphRAG follows with an overall score of 0.50, constrained by its reliance on entities and relationships extracted from the knowledge graph, which may be incomplete or missing, resulting in contextual gaps. Vector RAG performs best on easy questions (0.61) but its accuracy declines on medium and hard questions due to its dependence on direct semantic similarity, which limits effectiveness when relevant facts are not explicitly present in the retrieved chunks.
With respect to context relevance, GraphRAG consistently outperforms both Vector RAG and Hybrid GraphRAG across all question difficulty levels. Its use of structured entities and relationships enables the retrieval of concise, semantically aligned information while minimizing irrelevant content. In contrast, Hybrid GraphRAG yields the lowest context relevance (0.04). An example of this can be seen in Figure 4, where the integration of graph- and vector-based retrieval introduces additional verbosity, often including extraneous information that dilutes precision and reduces alignment with the query. In comparison, GraphRAG produces more direct and precise responses. Similarly, GraphRAG’s marginal lead in answer relevance (0.74) can also be attributed to its structured graph traversal, which yields more focused responses.

### v2

@@ Abstract
Generative AI (GenAI) is expected to play a pivotal role in enabling autonomous optimization in future wireless networks. Within the ORAN architecture, Large Language Models (LLMs) can be specialized to generate xApps and rApps by leveraging specifications and API definitions from the RAN Intelligent Controller (RIC) platform. However, fine-tuning base LLMs for telecom-specific tasks remains expensive and resource-intensive. Retrieval-Augmented Generation (RAG) offers a practical alternative through in-context learning, enabling domain adaptation without full retraining. While traditional RAG systems rely on vector-based retrieval, emerging variants such as GraphRAG and Hybrid GraphRAG incorporate knowledge graphs or dual retrieval strategies to support multi-hop reasoning and improve factual grounding. Despite their promise, these methods lack systematic, metric-driven evaluations, particularly in high-stakes domains such as ORAN. In this study, we conduct a comparative evaluation of Vector RAG, GraphRAG, and Hybrid GraphRAG using ORAN specifications. We assess performance across varying question complexities using established generation metrics: faithfulness, answer relevance, context relevance, and factual correctness. Results show that both GraphRAG and Hybrid GraphRAG outperform traditional RAG. Hybrid GraphRAG improves factual correctness by 8%, while GraphRAG improves context relevance by 11%.
@@ II Related Work
Figure 2: Overview of the experimental pipeline, illustrating the core components and data flow across the retrieval, generation and evaluation stages.
@@ III-C Hybrid GraphRAG
Figure 3: Example response generated by Vector RAG, GraphRAG, and Hybrid GraphRAG for a benchmark question from the ORAN-Bench-13K dataset.
@@ III-D Dataset
For this study, we utilized a corpus of 74 documents from the ORAN Alliance Specifications [28]. The evaluation was conducted using the ORAN-Bench-13K dataset [14], a benchmark specifically designed to assess the performance of LLMs within the ORAN context. The dataset categorizes questions into three levels of complexity: Easy, Intermediate, and Hard. We selected a stratified subset of 600 questions to ensure balanced coverage across all difficulty levels and representative topical diversity. The sampling procedure maintained proportional representation from categories such as network architecture, analytics and monitoring, anomaly detection, and protocol interpretation. Each question includes four answer options along with a ground truth label. The categorization reflects increasing levels of difficulty: (i) Easy: Questions targeting foundational concepts or factual knowledge (Simple QA) (ii) Intermediate: Questions requiring moderate reasoning, comprehension, or application of concepts (Complex Reasoning QA) (iii) Hard: Questions demanding deep understanding of ORAN standards and the ability to synthesize information across multiple documents (Multi-hop Reasoning)  [6].
@@ IV-B Results
Figure 4 visualizes the performance of each model across four core evaluation metrics stratified by question difficulty. To provide a broader perspective, Table II reports the mean of each evaluation metric aggregated across easy, medium, and hard questions. The corresponding standard deviations capture variability in performance across these difficulty levels. Each experiment is repeated thrice to ensure consistency and reliability. The results are discussed below, organized by evaluation metric.
TABLE II: Average Performance Across Question Difficulty Levels
@@
Faithfulness 0.59 $\pm$ 0.16 0.59 $\pm$ 0.16 0.55 $\pm$ 0.22
Factual Correctness 0.50 $\pm$ 0.17 0.58 $\pm$ 0.10 0.48 $\pm$ 0.12
Context Relevance 0.56 $\pm$ 0.10 0.45 $\pm$ 0.05 0.51 $\pm$ 0.11
Answer Relevance 0.74 $\pm$ 0.01 0.72 $\pm$ 0.01 0.73 $\pm$ 0.01
In terms of faithfulness, both GraphRAG and Hybrid GraphRAG outperform Vector RAG by 4%, with lower variability across difficulty levels. This suggests that the graph-based pipelines produce responses that are more consistently grounded in the retrieved context and are less susceptible to hallucinations.
In terms of factual correctness, Hybrid GraphRAG achieves the highest average score with low variability (0.58 ± 0.10) across all difficulty levels, as shown in Table II. Its performance remains stable due to its ability to compensate for limitations in individual retrieval strategies, falling back on vector-based retrieval when graph-derived context is insufficient, and vice versa. GraphRAG follows with an average score of 0.50, limited by its dependence on entities and relationships extracted from the knowledge graph, which may be incomplete or sparse. As shown in Figure 4, Vector RAG performs best on easy questions (0.61), but its accuracy drops on medium and hard questions due to its reliance on direct semantic similarity, which is less effective when relevant information is not explicitly retrieved.
With respect to context relevance, GraphRAG outperforms Hybrid GraphRAG by 11% on average across all question difficulty levels. This improvement is attributed to GraphRAG’s use of structured entities and relationships, which facilitates the retrieval of concise information while minimizing irrelevant content. In contrast, Hybrid GraphRAG exhibits the lowest context relevance score (0.45 ± 0.05), indicating consistently weaker alignment between retrieved context and query across difficulty levels. As shown in Figure 3, the response generated by Hybrid GraphRAG, though factually accurate, tends to include verbose or tangential details that dilute relevance and reduce semantic precision with respect to the query. Answer relevance remains consistent across models, with GraphRAG showing a slight lead (0.74 ± 0.01), likely due to its structured graph traversal that promotes more focused responses.